Liza's age is two more than thrice the age of her younger sister. If her younger sister is 5 years old, how old is Liza?
Answer by ikleyn(52781)

.

Liza's age = 3*5 + 2 = 15 + 2 = 17.      (*)


Liza is 17 years old.


The line (*) is the direct and literal translation of the problem' description.

Solved, answered and explained. And completed.